OpenStudy (anonymous):

28=8x + 12 - 7x show work

OpenStudy (anonymous):

The first step is to add 7x and 8x

OpenStudy (anonymous):

Then you subtract 12 on both sides of the equation

OpenStudy (anonymous):

So now you have 15x = 16

OpenStudy (anonymous):

now divide 15x/15 = 16/15 to both sides of the equation and you get x = 16/15

OpenStudy (anonymous):

Was that great Help or what?????

OpenStudy (anonymous):

Looks like all you did was solve it for her, so that was actually terrible help.